OpenStudy (anonymous):

how do you solve 2(x+3)-1=7x

OpenStudy (bradely):

2(x+3)-1=7x Using distributive property 2x+2(3)=7x 2x+6=7x 5x = 6 x = 6/5 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Thank you so much!